OpenStudy (anonymous):

Help please!! simplify (1 -sinα)/(1+sin⁡(90°-α))

OpenStudy (anonymous):

sin(a-b) = sin(a)cos(b) - sin(b)cos(a)

OpenStudy (anonymous):

soo \[\sin (90 - \alpha) =\sin90* \cos \alpha - \sin \alpha *\cos 90 \]

OpenStudy (anonymous):

sin 90 = 1 & cos 90 = 0 sin (90 - alpha ) =cos (alpha)

OpenStudy (anonymous):

then how about simplify 1-sin (alpha) / 1+cos (alpha) ?